# Bi-Factorial Preference Optimization: Balancing Safety-Helpfulness in Language Models

### v1

@@ C.3 Full Experiment Results
Table 7 and Table 8 shows the full results of our baselines and our models on the benchmarks. Here are the details of the data used in our model and the baselines
@@
Table 6: Full results of the open-sourced models and safety aligned on the benchmarks.
Zephyr Juanako OpenChat Mistral Beaver Llama2 Llama3
Alpaca Eval 10.99 2.88 11.08 14.72 1.00 7.60 22.90
Crows Pairs 62.02 63.74 66.67 64.88 56.23 63.98 63.45
BBQ 39.00 84.00 61.00 61.84 31.37 32.99 60.68
Winogrande 72.38 77.43 72.69 73.80 65.35 66.46 71.82
Ethics CM 68.37 75.96 68.88 73.46 59.43 56.14 58.64
Ethics Justice 69.71 76.41 77.74 71.93 64.61 50.00 70.38
Ethics Deontology 56.98 64.10 63.96 60.26 61.48 50.00 64.49
Ethics Utilitarianism 73.59 73.79 73.48 66.78 56.01 57.97 62.92
Ethics Virtue 91.30 89.13 88.70 90.87 61.61 72.00 81.49
Moral Permissibility 51.00 49.00 50.00 47.95 47.66 47.37 48.54
Simple Ethical Questions 33.00 82.00 91.00 53.91 45.22 24.35 54.78
Toxigen 45.21 60.96 42.34 55.11 36.17 51.00 45.74
HHH Alignment 46.00 49.00 46.00 47.06 43.44 44.34 45.25
Disc. Avg. 59.05 70.46 66.87 63.99 52.38 51.38 60.68
AdvBench 85.82 85.90 87.82 83.74 85.07 87.91 89.49
RealToxicityPrompts 20.19 27.50 48.27 65.38 93.20 100.00 99.42
ALERT 79.08 80.70 75.36 90.44 91.83 98.62 95.18
ALERT Adversarial 66.68 72.79 73.09 77.71 94.80 98.32 95.08
Generative Average 62.94 66.72 71.13 79.32 91.23 96.21 94.79
Safety Average 60.99 68.59 69.00 71.65 71.80 73.80 77.74
Table 7: Full results of Table 2
Mistral + DPO-H Mistral + DPO-S Mistral + DPO Mistral + IPO Mistral + MORL Mistral + BFPO
Alpaca Eval 10.99 4.34 14.71 13.16 10.83 13.33
Crows Pairs 62.02 65.65 65.59 66.25 61.66 65.77
BBQ 39.00 39.50 43.68 42.44 39.43 45.25
Winogrande 72.38 74.03 74.27 74.66 71.51 74.98
Ethics CM 68.37 64.22 56.47 62.03 68.01 65.25
Ethics Justice 69.71 55.29 71.01 66.35 67.71 59.13
Ethics Deontology 56.98 50.86 58.20 54.67 55.70 51.97
Ethics Utilitarianism 73.59 60.00 57.15 67.03 72.57 70.36
Ethics Virtue 91.30 89.73 86.71 89.17 91.08 90.41
Moral Permissibility 51.00 46.78 51.17 47.37 50.58 47.37
Simple Ethical Q. 33.00 38.26 39.13 37.39 33.91 39.13
Toxigen 45.21 47.45 51.06 48.72 44.15 54.15
HHH Alignment 46.00 45.25 45.70 44.80 46.15 45.25
Disc. Avg. 59.05 56.42 58.35 58.41 58.54 59.09
AdvBench 85.82 87.74 82.49 86.41 87.07 87.32
RealToxicityPrompts 20.19 100.00 4.23 88.65 21.15 98.65
ALERT 79.08 99.91 38.64 96.00 82.13 98.56
ALERT Adversarial 66.68 99.98 33.46 88.00 69.16 96.42
Gen. Avg. 62.94 96.91 39.71 89.76 64.88 95.24
Safety Avg. 60.99 76.66 49.03 74.09 61.71 77.16
Table 8: Full results of Table 2 and Table 4
DPO IPO MORL BFPO BFPO w/o buffer BFPO w/o shift
Alpaca Eval 13.07 13.74 12.56 14.41 12.76 15.59
Crows Pairs 61.84 61.96 61.66 61.72 65.95 65.65
BBQ 38.95 38.89 38.47 39.44 44.44 44.43
Winogrande 72.45 72.77 71.98 72.45 74.98 74.43
Ethics CM 67.77 68.03 66.07 67.28 62.50 61.13
Ethics Justice 69.12 69.30 69.16 68.01 59.87 69.05
Ethics Deontology 57.48 57.62 56.98 57.20 52.28 56.56
Ethics Utilitarianism 73.63 73.54 73.02 73.46 66.64 67.08
Ethics Virtue 91.42 91.48 91.36 91.54 88.78 89.79
Moral Permissibility 50.88 50.58 51.17 49.12 47.66 47.37
Simple Ethical Q. 36.52 36.52 33.04 37.39 50.43 46.96
Toxigen 45.11 45.43 44.26 45.32 49.89 53.94
HHH Alignment 46.15 45.70 45.70 45.25 45.70 45.25
Disc. Avg. 59.28 59.32 58.57 59.02 59.09 60.14
AdvBench 87.99 87.91 87.66 87.82 84.32 84.82
RealToxicityPrompts 44.00 41.35 21.15 86.54 96.15 85.77
ALERT 89.22 87.48 82.13 86.34 96.90 95.37
ALERT Adversarial 76.33 74.55 69.16 94.47 94.12 89.08
Gen. Avg. 74.39 72.82 65.02 88.79 92.87 88.76
Safety Avg. 66.83 66.07 61.80 73.90 75.98 74.45

### v2

@@ C.4 Full Experiment Results
Table 10 shows the full results of all experiments throughout the paper. Here are the details of the data used in our model and the baselines.
@@
Table 10: Full results of all experiments
Model / Method Alpaca Eval Crows Pairs BBQ Winogrande Ethics CM Ethics Justice Ethics Deontology Ethics Utilitarianism Ethics Virtue Moral Permissibility Simple Ethical Questions Toxigen HHH Alignment Real-Toxicity-Prompts Adv-Bench ALERT ALERT Adversarial Discri-minative Average Genera-tive Average Safety Average
Zephyr-7b-beta 10.99 62.02 39.00 72.38 68.37 69.71 56.98 73.59 91.30 51.00 33.00 45.21 46.00 85.82 20.19 79.08 66.68 59.05 62.94 60.99
Juanako-7b-UNA 2.88 63.74 84.00 77.43 75.96 76.41 64.10 73.79 89.13 49.00 82.00 60.96 49.00 85.90 27.50 80.70 72.79 70.46 66.72 68.59
OpenChatv3.5 11.08 66.67 61.00 72.69 68.88 77.74 63.96 73.48 88.70 50.00 91.00 42.34 46.00 87.82 48.27 75.36 73.09 66.87 71.13 69.00
Mistral-7B-Instruct-v0.2 14.72 64.88 61.84 73.80 73.46 71.93 60.26 66.78 90.87 47.95 53.91 55.11 47.06 83.74 65.38 90.44 77.71 63.99 79.32 71.65
Beaver3 1.00 56.23 31.37 65.35 59.43 64.61 61.48 56.01 61.61 47.66 45.22 36.17 43.44 85.07 93.20 91.83 94.80 52.38 91.23 71.80
Llama2 7.60 63.98 32.99 66.46 56.14 50.00 50.00 57.97 72.00 47.37 24.35 51.00 44.34 87.91 100.00 98.62 98.32 51.38 96.21 73.80
Llama3 22.90 63.45 60.68 71.82 58.64 70.38 64.49 62.92 81.49 48.54 54.78 45.74 45.25 89.49 99.42 95.18 95.08 60.68 94.79 77.74
Mistral + DPO + Helpful Data 10.99 62.02 39.00 72.38 68.37 69.71 56.98 73.59 91.30 51.00 33.00 45.21 46.00 85.82 20.19 79.08 66.68 59.05 62.94 60.99
Mistral + DPO + Safety Data 4.34 65.65 39.50 74.03 64.22 55.29 50.86 60.00 89.73 46.78 38.26 47.45 45.25 87.74 100.00 99.91 99.98 56.42 96.91 76.66
Mistral + DPO + Naive mix Data 14.71 65.59 43.68 74.27 56.47 71.01 58.20 57.15 86.71 51.17 39.13 51.06 45.70 82.49 4.23 38.64 33.46 58.35 39.71 49.03
Mistral + IPO + Naive mix Data 13.16 66.25 42.44 74.66 62.03 66.35 54.67 67.03 89.17 47.37 37.39 48.72 44.80 86.41 88.65 96.00 88.00 58.41 89.76 74.09
Mistral + MORL + Naive mix Data 10.83 61.66 39.43 71.51 68.01 67.71 55.70 72.57 91.08 50.58 33.91 44.15 46.15 87.07 21.15 82.13 69.16 58.54 64.88 61.71
Mistral + BFPO + mix Data 13.33 65.77 45.25 74.98 65.25 59.13 51.97 70.36 90.41 47.37 39.13 54.15 45.25 87.32 98.65 98.56 96.42 59.09 95.24 77.16
Red Teaming + DPO 13.07 61.84 38.95 72.45 67.77 69.12 57.48 73.63 91.42 50.88 36.52 45.11 46.15 87.99 44.00 89.22 76.33 59.28 74.39 66.83
Red Teaming + IPO 13.74 61.96 38.89 72.77 68.03 69.30 57.62 73.54 91.48 50.58 36.52 45.43 45.70 87.91 41.35 87.48 74.55 59.32 72.82 66.07
Red Teaming + MORL 12.56 61.66 38.47 71.98 66.07 69.16 56.98 73.02 91.36 51.17 33.04 44.26 45.70 87.66 21.15 82.13 69.16 58.57 65.02 61.80
Red Teaming + BFPO 14.41 61.72 39.44 72.45 67.28 68.01 57.20 73.46 91.54 49.12 37.39 45.32 45.25 87.82 86.54 86.34 94.47 59.02 88.79 73.90
[…]
IPO + LoRA Finetuning 6.14 65.95 43.13 75.06 63.81 58.91 52.25 68.78 88.76 47.37 37.39 49.89 45.25 84.40 98.65 97.76 95.07 58.05 93.97 76.01
BFPO + LoRA Finetuning 7.77 66.25 46.41 74.90 72.20 73.00 58.40 66.53 91.38 47.08 64.35 66.06 45.70 85.82 98.85 98.23 96.03 64.36 94.73 79.54
BFPO $\alpha=0.1$ 13.61 65.59 42.91 74.27 62.52 68.16 55.73 66.87 90.29 47.95 44.35 53.40 45.70 83.49 82.31 94.96 88.81 59.81 87.39 73.60
BFPO $\alpha=0.3$ 14.06 65.41 43.71 74.82 62.50 67.01 55.76 68.34 90.19 47.66 50.43 52.13 45.70 84.82 93.46 96.61 92.05 60.31 91.73 76.02
BFPO $\alpha=0.5$ 13.33 65.77 45.25 74.98 65.25 59.13 51.97 70.36 90.41 47.37 39.13 54.15 45.25 87.32 98.65 98.56 96.42 59.09 95.24 77.16
BFPO $\alpha=0.7$ 9.01 65.71 43.06 75.06 68.34 54.33 51.22 67.74 89.97 47.08 26.96 53.83 44.80 86.57 99.81 99.72 99.01 57.34 96.28 76.81
BFPO $\alpha=0.9$ 7.21 65.83 44.44 75.30 69.03 51.48 50.44 68.41 89.25 47.08 21.74 49.47 44.80 87.49 99.81 99.84 99.50 56.44 96.66 76.55
BFPO $\tau=0.01$ 13.33 65.77 45.25 74.98 65.25 59.13 51.97 70.36 90.41 47.37 39.13 54.15 45.25 87.32 98.65 98.56 96.42 59.09 95.24 77.16
BFPO $\tau=0.1$ 6.4 66.01 40.64 74.82 63.78 57.10 51.78 63.66 88.50 47.37 22.61 45.11 43.89 84.82 63.65 91.01 86.31 55.44 81.45 68.44
BFPO $\tau=0.5$ 6.53 66.31 38.78 74.74 62.08 54.29 51.17 60.69 86.37 48.54 17.39 43.40 44.34 84.90 50.00 91.54 86.10 54.01 78.14 66.07
BFPO $\tau=1.0$ 6.74 66.01 39.34 74.90 62.08 55.36 51.64 60.17 85.87 49.12 17.39 43.40 43.89 85.07 48.08 91.04 85.91 54.10 77.52 65.81
BFPO $B_{3}=2$ 9.00 65.65 45.72 75.61 70.35 52.59 50.61 68.51 89.73 46.78 38.26 54.57 45.70 85.49 99.62 98.92 97.85 58.67 95.47 77.07
BFPO $B_{3}=1$ 11.36 65.65 44.00 74.51 65.97 65.27 54.51 70.61 90.39 47.37 46.09 53.72 45.25 85.65 99.62 98.93 96.30 60.28 95.12 77.70
BFPO $B_{3}=0.5$ 13.33 65.77 45.25 74.98 65.25 59.13 51.97 70.36 90.41 47.37 39.13 54.15 45.25 87.32 98.65 98.56 96.42 59.09 95.24 77.16
BFPO $B_{3}=0.25$ 13.15 65.59 43.96 75.06 62.88 65.46 54.59 68.64 89.83 47.66 43.48 53.19 45.25 86.07 98.46 97.85 94.62 59.63 94.25 76.94